# Non-Adaptive Multi-Stage Algorithm and Bounds for Group Testing with Prior Statistics

### v3

@@ A-B Definite Defectives (DD)
A-C List Viterbi Algorithm (LVA)
@@ Appendix B Probability of Error (Proof of Lemma 1)
$=\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)\cdot$
$\hskip 14.22636pt\frac{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}P\left(\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}P_{W}\left(\omega^{\prime}\right)^{s}}{P\left(\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}P_{W}\left(\omega^{*}\right)^{s}}$
@@
$=\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}}{P_{W}\left(\omega^{*}\right)^{s}}\cdot$
$\hskip 71.13188pt\frac{\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}}$
$\overset{(a)}{\leq}\Biggl{[}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)}{P_{W}\left(\omega^{*}\right)^{s}}\cdot$
$\hskip 71.13188pt\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}}\Biggr{]}^{\rho}$
$=\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{s\rho}}\cdot$
$\hskip 71.13188pt\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s\rho}},$
@@
$=\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(\mathbf{X}_{TP},Y^{T}\lvert\mathbf{X}_{FN}\right)P\left(\mathbf{X}_{FN}\right)\cdot$
$\hskip 22.76228ptP\left(E_{i}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$
$\overset{(a)}{\leq}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{s\rho}}\cdot$
$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)\cdot$
$\hskip 39.83368ptP\left(\mathbf{X}_{FN}\right)\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s\rho}}$
$\overset{(b)}{\leq}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right))}}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{\frac{\rho}{\left(1+\rho\right)}}}\cdot$
$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)P\left(\mathbf{X}_{FN}\right)\cdot$
$\hskip 45.52458pt\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right))^{\frac{\rho}{1+\rho}}}$
$=\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{\frac{\rho}{\left(1+\rho\right)}}}\cdot$
$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
$=\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\cdot$
$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
$=\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\cdot$
$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},$
@@
$\leq\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}$
$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
$=\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)^{1-\frac{\rho}{1+\rho}}{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}$
$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
$=\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)^{\frac{1}{1+\rho}}{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}$
$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
=$\overset{(a)}{=}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{1+\rho}}\right]$
$\hskip 22.76228pt{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}$
$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
$\overset{(b)}{\leq}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}{\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{1+\rho}}$
$\hskip 22.76228pt\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
=$\overset{(c)}{=}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}{\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{1+\rho}}$
$\hskip 19.91684pt\left[\sum_{Y^{T}}\sum_{X_{TP}}\left(\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y^{T},X_{TP}\lvert X_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}\right]^{T}$
@@
where (a) follows from Lemma 2 with $\alpha=\frac{1}{1+\rho}$, (b) follows from the fact that for a fixed $S_{1}\in\mathcal{T}\mathcal{P}_{i}$, ~$\tilde{\mathcal{I}}^{j}_{dif}\subset\left\{\omega=\left(S_{1},S_{2}\right):S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)\right\}$, (c) follows since $X_{FP},X_{TP}$ represent a single column of $\mathbf{X}_{FP},\mathbf{X}_{TP}$ respectively, as the elements of $\mathbf{X}$ are drawn in i.i.d. fashion, and (d) follows for
$\begin{split}&E_{0}(\rho)=\\ &-\log\sum_{Y^{T}}\sum_{X_{TP}}\left(\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y^{T},X_{TP}\lvert X_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},\end{split}$

### v4

@@ A-B Definite Defectives (DD)
A-C Priors Update
The algorithm translates the information gathered during the first stage of MSGT into the state space. In particular, if an element is detected as defective, its probability of being defective is updated to $1$. Additionally, the transition probabilities are modified so that the transition to the defective state becomes deterministic.
Algorithm 4 updatePriors
Input: $\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N},\mathcal{P}^{(S_{1,2})},\mathcal{P}^{(DD)}$
Output: $\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}$
1:for each $i\in\mathcal{P}^{(\mathrm{DD})}$ do
2:    $\pi_{i}\leftarrow 0$
3:    if $i\neq 1$ then
4:         $\boldsymbol{\Phi}_{i-1}[:,0]\leftarrow\boldsymbol{0}$
5:         $\boldsymbol{\Phi}_{i-1}[:,1]\leftarrow\boldsymbol{1}$
6:    end if
7:end for
8:return $\{\mathbf{\pi}_{i},\mathbf{\Phi}_{i}\}_{i=1}^{N}$
A-D List Viterbi Algorithm (LVA)
@@ Appendix B Probability of Error (Proof of Lemma 1)
$=\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)\cdot\frac{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}P\left(\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}P_{W}\left(\omega^{\prime}\right)^{s}}{P\left(\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}P_{W}\left(\omega^{*}\right)^{s}}$
@@
$=\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}}{P_{W}\left(\omega^{*}\right)^{s}}\cdot\frac{\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}}$
$\overset{(a)}{\leq}\left[\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)}{P_{W}\left(\omega^{*}\right)^{s}}\cdot\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}}\right]^{\rho}$
$=\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{s\rho}}\cdot\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s\rho}},$
@@
$=\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(\mathbf{X}_{TP},Y^{T}\lvert\mathbf{X}_{FN}\right)P\left(\mathbf{X}_{FN}\right)\cdot P\left(E_{i}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$
$\overset{(a)}{\leq}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{s\rho}}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)\cdot P\left(\mathbf{X}_{FN}\right)\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s\rho}}$
$\overset{(b)}{\leq}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right))}}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{\frac{\rho}{\left(1+\rho\right)}}}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)P\left(\mathbf{X}_{FN}\right)\cdot\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right))^{\frac{\rho}{1+\rho}}}$
$=\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{\frac{\rho}{\left(1+\rho\right)}}}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
$=\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
$=\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},$
@@
$\leq\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
$=\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)^{1-\frac{\rho}{1+\rho}}{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
$=\sum_{\omega^{*}\in\left[\binom{N}{K}\right]}P_{W}\left(\omega^{*}\right)^{\frac{1}{1+\rho}}{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
=$\overset{(a)}{=}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{1+\rho}}\right]{\left[\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{\rho}}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
$\overset{(b)}{\leq}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}{\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{1+\rho}}\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
=$\overset{(c)}{=}\frac{1}{\binom{K}{i}}\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}{\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{1+\rho}}\left[\sum_{Y^{T}}\sum_{X_{TP}}\left(\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y^{T},X_{TP}\lvert X_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}\right]^{T}$
@@
where (a) follows from Lemma 2 with $\alpha=\frac{1}{1+\rho}$, (b) follows from the fact that for a fixed
$S_{1}\in\mathcal{T}\mathcal{P}_{i},\quad\tilde{\mathcal{I}}^{j}_{dif}\subset\left\{\omega=\left(S_{1},S_{2}\right):S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)\right\},$
(c) follows since $X_{FP},X_{TP}$ represent a single column of $\mathbf{X}_{FP},\mathbf{X}_{TP}$ respectively, as the elements of $\mathbf{X}$ are drawn in i.i.d. fashion, and (d) follows for
$\begin{split}&E_{0}(\rho)=-\log\sum_{Y^{T}}\sum_{X_{TP}}\left(\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y^{T},X_{TP}\lvert X_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},\end{split}$